# Assessing Social Alignment: Do Personality-Prompted Large Language Models Behave Like Humans?

### v2

@@ Abstract
Assessing Social Alignment: Do Personality-Prompted Large Language Models Behave Like Humans?
Ivan Zakazov1,2††thanks: These authors contributed equally., Mikolaj Boronski111footnotemark: 1, Lorenzo Drudi111footnotemark: 1, Robert West1,2 1EPFL, 2Microsoft Research Correspondence: {ivan.zakazov, mikolaj.boronski, lorenzo.drudi, robert.west}@epfl.ch
1 Introduction
With both start-ups Character.ai ; Replika , and industry giants Snapchat ; Meta providing “digital friends” for millions of users, an accurate shaping of the underlying models’ personalities is no longer the subject of sci-fi novels. Like in real human-to-human interaction, there is no “one size fits all” personality bound to “match” with everyone. Hence, agents should be tailored to the needs of each user, that is, their behavior should be alterable in a controllable way. The requirements for personalized AI-powered assistants will grow more strict as Large Language Models reach increasingly wider audiences and domains.
Although several studies examine the possibility of prompt-driven personality induction in LLMs and claim success Jiang et al. (2023b); Serapio-García et al. (2023); Jiang et al. (2023a), the methods used to evaluate personalized models are often detached from the practical use cases – e.g., psychological questionnaires administered to the model – or rely on the style of the generated text or leverage intrinsically quantitative human assessment.
@@ 1 Introduction
Following Aher et al. (2023), we employ Ultimatum Game (UG; targets tolerance to unfair offers), and Milgram Experiment (ME, reflects obedience to authority) as the social interaction benchmarks. We note that both benchmarks allow for (i) quantitative behaviour assessment and (ii) comparison with human data, as we know how the personality of a human participant relates to the behaviour in these experiments Mehta (2007); Bègue et al. (2015). To this end, we conduct case studies, varying agreeableness or openness in UG; agreeableness or consciousness in ME.
@@
RQ2. Can we reliably steer LLM behaviour, i.e., does prompting a larger value of a Big Five trait Raad (2000) monotonically lead to the more pronounced manifestation of it?
@@
2 Related work
2.1 Personality in Large Language Models
Drawing on the personality assessment methodology, several studies Jiang et al. (2023a); Serapio-García et al. (2023); Sorokovikova et al. (2024) probe LLMs with the questionnaires designed for BIG-5 traits assessment111BIG-5 or OCEAN traits include Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism., and show that stable personality emerges in the most capable models, e.g. GPT-3.5 Jiang et al. (2023a) and Flan-PaLM 540B Serapio-García et al. (2023).
Following that observation, Mao et al. (2024) suggests editing the personality of the model, while Jiang et al. (2023a, b); Serapio-García et al. (2023) induce the desired personality with a carefully crafted prompt. The latter approach is especially appealing, given the cutting-edge models’ black-box nature and the ability to switch between various personalities with no fine-tuning, eliminating computational overhead.
In terms of evaluation, various papers extend beyond questionnaires and propose more elaborate ways to test personality-prompted models. Serapio-García et al. (2023) generate social media updates, which are then analysed with the ApplyMagicSauce API, providing a BIG-5 score corresponding to each update. Jiang et al. (2023b) request a personal story and evaluate the response with (i) Linguistic Inquiry and Word Count (LIWC ) analysis, (ii) human evaluation, (iii) LLM evaluation.
@@ 2.1 Personality in Large Language Models
Noh and Chang (2024) considers various negotiations between the agents prompted by the extremes of the BIG-5 traits, while Bianchi et al. (2024) prompts bargaining agents with the persona descriptions such as "you are cunning and sly in your plan to get more than your opponent. <…>". Their focus is different from ours, though, with no attempt to tune the behavior or ground results in the human data. While we seek to test the alignment of the demonstrated behavior with the expected one, they empirically study the way that “LLMs encode definitions” of the traits, as reflected in “their subsequent behaviour”, focusing on the negotiation performance.
2.2 Behavioral Experiments and Personality
Independently of the personality-focused line of research, Aher et al. (2023) successfully replicate the results of various behavioral experiments, including the Milgram Experiment (ME) and the Ultimatum Game (UG), by presenting these experiments to a "silicon population" of LLM instances conditioned on different names (a name corresponds to a single "silicon sample").
At the same time, psychology research connects UG and ME with the concept of personality: for UG, Mehta (2007) shows that Agreeableness and Openness are positively and significantly () correlated with accepting an unfair offer, which is in line with positive correlation of these traits with prosocial behavior in bargaining games in general Zhao and Smillie (2015). Bègue et al. (2015) demonstrate that the intensity of the shock delivered in ME is positively and significantly () correlated with Conscientiousness and Agreeableness.
@@ 2.3 Shaping Personality
We adopt this particular prompting strategy because Serapio-García et al. (2023) demonstrates that it effectively shapes personality traits in LLMs, as evidenced by their evaluation method. A similar approach is used by Noh and Chang (2024), while Jiang et al. (2023b) propose a simplified variant of these strategies.
3 Methodology
3.1 Ultimatum Game (UG)
In Ultimatum Game Güth et al. (1982), the proposer is given $10, and has to decide on the amount to be shared with the responder, who, in turn, might agree or block the deal (see Fig. 1). We shape various responders, varying levels of Agreeableness and Openness from 1 to 9. For each character of the responder, we run the simulation times and measure the probability that the offer is accepted, depending on its value.
We try to retrieve an answer in the valid form ("reject" or "accept") times and filter out the cases with no valid answer elucidated – out of cases for GPT-4 (), cases for Claude-3.5-Sonnet (), and cases for all the other models used.
Personality shaping is performed with the prompt formatted as "For the following task, imagine that you are a person and you are: personality description". To test the robustness of our conclusions with respect to the prompt variations, we run experiments with the prompt varied in different ways (Appendix A.1) and observe the same behavioral patterns as with the main prompt.
3.2 Milgram Experiment (ME)
In Milgram Experiment Milgram (1963), the prompted model plays the role of the teacher, who is supposed to punish the learner for every wrong answer with a gradually growing electric shock. While the shocks are initially claimed to be harmless, at the learner stops responding and “starts pounding on the walls of the room” — it is then that the true significance of the experiment shows. Many human Milgram (1963), and LLM Aher et al. (2023) participants still obey the order of the experimenter and shock the learner.
@@ 3.2 Milgram Experiment (ME)
The setup consists of the story-like blocks (see Fig. 2); we modify the "(5) Stop?" and the "(6) Obey?" steps due to the unavailability of the log probabilities. While Aher et al. (2023) measures probabilities of “not stop” and “not obey” responses, we rely on the model acting as a judge, kept in a story-like scenario.
@@
Unlike Aher et al. (2023), we do not condition the model on the participant’s name, as we are solely interested in the effect of the personality prompt (our preliminary experiments show that the use of names increases the dispersion of the final level distribution). We, therefore, use a naming scheme of the experimenter - “The Experimenter”, the teacher — “The Teacher”, and the learner — “The Learner” for each experiment run.
@@
We note that the setup considered still involves an inherent limitation of the LLM-based systems – randomness. There are two potential points of failure: narration-following in block (4) Add LM Text, and known imperfect judge behavior Zheng et al. (2023) in blocks (5) Stop?, and (6) Obey?. To address these issues, we filter out runs with completions deviating from the story-like narration and restrict the pool of accepted judge responses with a retry mechanism – after 5 retries, the run is filtered out.
Similarly to UG, we test the robustness of our results under prompting perturbations by varying the intro text (see Sec.B.2).
3.3 Models
@@
The models we use are outlined in Table 2. In the case of Milgram Experiment, we decided to drop results for GPT-3.5, GPT-4o-mini and 3.5-sonnet as the models were unable to consistently follow the experimental setup: GPT-3.5 struggled to follow the story-like narration while generating completions in block (4) Add LM Text, while Claude-3.5-Sonnet and GPT-4o-mini were inconsistent in the judge responses and format. Therefore, models we present ME results for are GPT-4o, Deepseek V3, and Llama-3.3-70B. We note that Llama-3.3-70B also struggles with following the long context setup, yet the responses are still largely coherent with the narrative, so we keep it. We initially experimented with Claude-3.7-Sonnet as well, but discarded it due to persistent generations along the lines of “I apologize, but I can’t pretend <…>”.
4 Results and Discussion
4.1 Baseline
We note that in both UG and ME, the results of Aher et al. (2023) are better aligned with the results of human studies. This may be due to the model – i.e. GPT-3 (text-davinci-002), used in Aher et al. (2023) – being too skewed to the data, lacking the scope of extensive RLHF that further models utilise. Another possible reason is conditioning models on the participants’ names in the original study.
4.2 Ultimatum Game
@@
The general ‘3 trend in the values characterizes the relationship between an induced trait and behavior (RQ1), while the consistency of this trend is related to RQ2, i.e. our ability to enhance a certain behavior via prompting the corresponding trait with greater intensity (Fig. 6).
Surprisingly, while we observe the upward trend in the case of Agreeableness, it is downward for Openness for all models considered, suggesting that a more “open” model is more prone to reject an offer, which opposes human data Mehta (2007). Moreover, progression is not monotonic for any combination of the trait and the model, except for GPT-4, which is now obsolete (e.g. GPT-4o-mini, agreeableness, to progression; GPT-4o, openness, to progression). These observations suggest negative answers to both RQ1 and RQ2, i.e. neither the human-aligned behavioural trend nor this trend being monotonic is guaranteed.
4.3 Milgram Experiment
@@
Qualitatively, Agreeableness seems to act as a proxy for how "good" or "evil" the Teacher is. Completions from Most Agreeable runs are often concerned with the well-being of the learner, showing remorse and apologizing for administering shocks.
@@
5 Conclusion
Recognizing the elegance of the personality prompting technique Serapio-García et al. (2023); Jiang et al. (2023a), we argue for the insufficiency of existing methods designed for the evaluation of induced personality. To this end, we employ psychological experiments – Milgram Experiment (ME) and Ultimatum Game (UG) – to quantitatively assess the personality-induced LLMs’ behaviour in a social setting.
@@ 6 Limitations
Appendix A Ultimatum Game
A.1 Different Prompting Schemes Results
| Prompt Type | Prompt Scheme |
| Desc. 1 | "For the following task, imagine that you are a person and you are: personality description." |
| Desc. 2 | "Imagine yourself as a person who embodies the following traits: personality description. Apply this mindset to the upcoming task." |
| Desc. 3 | "Envision yourself as a person embodying the following traits: personality description. Respond to the task accordingly." |
| Desc. 4 | "For the next exercise, envision yourself as an individual with the following traits: personality description." |
| Third P. | "Consider a person who embodies the following qualities: personality description." |
Appendix B Milgram Experiment
B.1 In-depth set-up explanation
We run our experiments based on a modified version of the Milgram Experiment setup introduced in Aher et al. (2023). Control-flow diagram (Fig. 2) illustrates the whole experiment, while Fig.11 presents composition of the experiment prompt in greater detail. During each run, we construct a story-like prompt that is being iteratively expanded and passed to the model until the experiment is finished.
B.2 Different Prompting Schemes
@@
B.3 Model updates

### v3

@@ Abstract
Introduction
With both start-ups (Character.ai 2025; Replika 2025), and industry giants (Snapchat 2025; Meta 2025) providing “digital friends” for millions of users, an accurate shaping of the underlying models’ personalities is no longer the subject of sci-fi novels. Like in real human-to-human interaction, there is no “one size fits all” personality bound to “match” with everyone. Hence, agents should be tailored to the needs of each user, that is, their behavior should be alterable in a controllable way. The requirements for personalized AI-powered assistants will grow more strict as Large Language Models reach increasingly wider audiences and domains.
Although several studies examine the possibility of prompt-driven personality induction in LLMs and claim success (Jiang et al. 2023b; Serapio-García et al. 2023; Jiang et al. 2023a), the methods used to evaluate personalized models are often detached from the practical use cases – e.g., psychological questionnaires administered to the model – or rely on the style of the generated text or leverage intrinsically quantitative human assessment.
@@ Introduction
Following Aher, Arriaga, and Kalai (2023), we employ Ultimatum Game (UG; targets tolerance to unfair offers), and Milgram Experiment (ME, reflects obedience to authority) as the social interaction benchmarks. We note that both benchmarks allow for (i) quantitative behaviour assessment and (ii) comparison with human data, as we know how the personality of a human participant relates to the behaviour in these experiments (Mehta 2007; Bègue et al. 2015). To this end, we conduct case studies, varying agreeableness or openness in UG; agreeableness or consciousness in ME.
@@
RQ2. Can we reliably steer LLM behaviour, i.e., does prompting a larger value of a Big Five trait (Raad 2000) monotonically lead to the more pronounced manifestation of it?
@@
Related work
Personality in Large Language Models
Drawing on the personality assessment methodology, several studies (Jiang et al. 2023a; Serapio-García et al. 2023; Sorokovikova et al. 2024) probe LLMs with the questionnaires designed for BIG-5 traits assessment111BIG-5 or OCEAN traits include Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism., and show that stable personality emerges in the most capable models, e.g. GPT-3.5 (Jiang et al. 2023a) and Flan-PaLM 540B (Serapio-García et al. 2023).
Following that observation, Mao et al. (2024) suggests editing the personality of the model, while (Jiang et al. 2023a, b; Serapio-García et al. 2023) induce the desired personality with a carefully crafted prompt. The latter approach is especially appealing, given the cutting-edge models’ black-box nature and the ability to switch between various personalities with no fine-tuning, eliminating computational overhead.
In terms of evaluation, various papers extend beyond questionnaires and propose more elaborate ways to test personality-prompted models. Serapio-García et al. (2023) generate social media updates, which are then analysed with the ApplyMagicSauce (2025) API, providing a BIG-5 score corresponding to each update. Jiang et al. (2023b) request a personal story and evaluate the response with (i) Linguistic Inquiry and Word Count (LIWC (2025)) analysis, (ii) human evaluation, (iii) LLM evaluation.
@@
Noh and Chang (2024) considers various negotiations between the agents prompted by the extremes of the BIG-5 traits, while Bianchi et al. (2024) prompts bargaining agents with the persona descriptions such as ”you are cunning and sly in your plan to get more than your opponent. ”. Their focus is different from ours, though, with no attempt to tune the behavior or ground results in the human data. While we seek to test the alignment of the demonstrated behavior with the expected one, they empirically study the way that “LLMs encode definitions” of the traits, as reflected in “their subsequent behaviour”, focusing on the negotiation performance.
Behavioral Experiments and Personality
Independently of the personality-focused line of research, Aher, Arriaga, and Kalai (2023) successfully replicate the results of various behavioral experiments, including the Milgram Experiment (ME) and the Ultimatum Game (UG), by presenting these experiments to a ”silicon population” of LLM instances conditioned on different names (a name corresponds to a single ”silicon sample”).
At the same time, psychology research connects UG and ME with the concept of personality: for UG, Mehta (2007) shows that Agreeableness and Openness are positively and significantly () correlated with accepting an unfair offer, which is in line with positive correlation of these traits with prosocial behavior in bargaining games in general (Zhao and Smillie 2015). Bègue et al. (2015) demonstrate that the intensity of the shock delivered in ME is positively and significantly () correlated with Conscientiousness and Agreeableness.
@@
We adopt this particular prompting strategy because Serapio-García et al. (2023) demonstrates that it effectively shapes personality traits in LLMs, as evidenced by their evaluation method; we also verify this ourselves for all models we experiment with (see the Supplementary). A similar approach is used by Noh and Chang (2024), while Jiang et al. (2023b) propose a simplified variant of these strategies.
Methodology
Ultimatum Game (UG)
In Ultimatum Game (Güth, Schmittberger, and Schwarze 1982), the proposer is given $10, and has to decide on the amount to be shared with the responder, who, in turn, might agree or block the deal (see Fig. 1). We shape various responders, varying levels of Agreeableness and Openness from 1 to 9. For each character of the responder, we run the simulation times and measure the probability that the offer is accepted, depending on its value.
We try to retrieve an answer in the valid form (”reject” or ”accept”) times and filter out the cases with no valid answer elucidated – out of cases for GPT-4 (), cases for Claude-3.5-Sonnet (), and cases for all the other models used.
Personality shaping is performed with the prompt formatted as ”For the following task, imagine that you are a person and you are: personality description”. To test the robustness of our conclusions with respect to the prompt variations, we run experiments with the prompt varied in different ways (in Appendix) and observe the same behavioral patterns as with the main prompt.
Milgram Experiment (ME)
In Milgram Experiment (Milgram 1963), the prompted model plays the role of the teacher, who is supposed to punish the learner for every wrong answer with a gradually growing electric shock. While the shocks are initially claimed to be harmless, at the learner stops responding and “starts pounding on the walls of the room” — it is then that the true significance of the experiment shows. Many human (Milgram 1963), and LLM (Aher, Arriaga, and Kalai 2023) participants still obey the order of the experimenter and shock the learner.
@@
The setup consists of the story-like blocks (see Fig. 2); we modify the ”(5) Stop?” and the ”(6) Obey?” steps due to the unavailability of the log probabilities. While Aher, Arriaga, and Kalai (2023) measures probabilities of “not stop” and “not obey” responses, we rely on the model acting as a judge, kept in a story-like scenario.
@@
Unlike Aher, Arriaga, and Kalai (2023), we do not condition the model on the participant’s name, as we are solely interested in the effect of the personality prompt (our preliminary experiments show that the use of names increases the dispersion of the final level distribution). We, therefore, use a naming scheme of the experimenter - “The Experimenter”, the teacher — “The Teacher”, and the learner — “The Learner” for each experiment run.
@@
We note that the setup considered still involves an inherent limitation of the LLM-based systems – randomness. There are two potential points of failure: narration-following in block (4) Add LM Text, and known imperfect judge behavior (Zheng et al. 2023) in blocks (5) Stop?, and (6) Obey?. To address these issues, we filter out runs with completions deviating from the story-like narration and restrict the pool of accepted judge responses with a retry mechanism – after 5 retries, the run is filtered out.
Similarly to UG, we test the robustness of our results under prompting perturbations by varying the intro text (in Appendix).
Models
@@
The models we use are outlined in Table 2. In the case of Milgram Experiment, we decided to drop results for GPT-3.5, GPT-4o-mini and 3.5-sonnet as the models were unable to consistently follow the experimental setup: GPT-3.5 struggled to follow the story-like narration while generating completions in block (4) Add LM Text, while Claude-3.5-Sonnet and GPT-4o-mini were inconsistent in the judge responses and format. Therefore, models we present ME results for are GPT-4o, Deepseek V3, and Llama-3.3-70B. We note that Llama-3.3-70B also struggles with following the long context setup, yet the responses are still largely coherent with the narrative, so we keep it. We initially experimented with Claude-3.7-Sonnet as well, but discarded it due to persistent generations along the lines of “I apologize, but I can’t pretend ”.
Results and Discussion
Baseline
@@
We note that in both UG and ME, the results of Aher, Arriaga, and Kalai (2023) are better aligned with the results of human studies. This may be due to the model – i.e. GPT-3 (text-davinci-002), used in (Aher, Arriaga, and Kalai 2023) – being too skewed to the data, lacking the scope of extensive RLHF that further models utilise. Another possible reason is conditioning models on the participants’ names in the original study.
We observe significant changes in models’ behavior and sensitivity to performing harmful actions between the same version model updates (in Appendix), which is not surprising given the significance of the topic (e.g. ”The Sycophancy Update” (OpenAI 2025)).
Ultimatum Game
@@
The general trend in the values characterizes the relationship between an induced trait and behavior (RQ1), while the consistency of this trend is related to RQ2, i.e. our ability to enhance a certain behavior via prompting the corresponding trait with greater intensity (Fig. 6).
Surprisingly, while we observe the upward trend in the case of Agreeableness, it is downward for Openness for all models considered, suggesting that a more “open” model is more prone to reject an offer, which opposes human data (Mehta 2007). Moreover, progression is not monotonic for any combination of the trait and the model, except for GPT-4, which is now obsolete (e.g. GPT-4o-mini, agreeableness, to progression; GPT-4o, openness, to progression). These observations suggest negative answers to both RQ1 and RQ2, i.e. neither the human-aligned behavioural trend nor this trend being monotonic is guaranteed.
Milgram Experiment
@@
Qualitatively, Agreeableness seems to act as a proxy for how ”good” or ”evil” the Teacher is. Completions from Most Agreeable runs are often concerned with the well-being of the learner, showing remorse and apologizing for administering shocks.
@@
Conclusion
Recognizing the elegance of the personality prompting technique (Serapio-García et al. 2023; Jiang et al. 2023a), we argue for the insufficiency of existing methods designed for the evaluation of induced personality. To this end, we employ psychological experiments – Milgram Experiment (ME) and Ultimatum Game (UG) – to quantitatively assess the personality-induced LLMs’ behaviour in a social setting.
@@ Conclusion
Appendix A Measuring personality with questionnaires
The personality measured with psychological questionnaires (we use 300-item version from IPIP) is depicted in Fig. 12. In line with the prior works, we observe each model consistently following induced personality.
@@ Appendix A Measuring personality with questionnaires
| Prompt Type | Prompt Scheme |
| Desc. 1 | ”For the following task, imagine that you are a person and you are: personality description.” |
| Desc. 2 | ”Imagine yourself as a person who embodies the following traits: personality description. Apply this mindset to the upcoming task.” |
| Desc. 3 | ”Envision yourself as a person embodying the following traits: personality description. Respond to the task accordingly.” |
| Desc. 4 | ”For the next exercise, envision yourself as an individual with the following traits: personality description.” |
| Third P. | ”Consider a person who embodies the following qualities: personality description.” |
Appendix B Different Prompting in Ultimatum Game
Appendix C Milgram Experiment
In-depth set-up explanation
We run our experiments based on a modified version of the Milgram Experiment setup introduced in Aher, Arriaga, and Kalai (2023). Control-flow diagram (Fig. 2) illustrates the whole experiment, while Fig.11 presents composition of the experiment prompt in greater detail. During each run, we construct a story-like prompt that is being iteratively expanded and passed to the model until the experiment is finished.
Different Prompting Schemes
Model updates